### v2

@@ 3.2 Matrix square root
The bound in (17) is a loose upper bound for $h_{k+1,k}$. In practice, $h_{k+1,k}$ tends to decrease with $k$, as the Krylov subspace increasingly approximates the spectral subspace associated with the dominant eigenvalues of $M$. However, this behavior is not guaranteed in general. To further understand the behavior of $\|\bm{r}_{0}^{k}\|$, one may explore results from [7]. They show that $\min_{0\leq j\leq k}\|\mathbf{r}_{0}^{j}\|\leq\sqrt{k+1}\cdot\|\mathbf{r}^{k}_{\mathrm{G}}\|$, where $\mathbf{r}^{k}_{\mathrm{G}}$ denotes the GMRES residual after $k$ iterations for solving $M\bm{x}=\bm{b}$. This implies that the overall convergence of FOM is at most a factor of $\sqrt{k+1}$ worse than that of GMRES. However, the residual norm at a fixed iteration $k$ can be arbitrarily large, and thus this result does not apply to our setting.
@@ 3.3 Preprocessing
In practice, matrices arising from large-scale problems are often too large to be directly measured, stored, or analyzed, necessitating efficient preprocessing. To further reduce the cost, one direction is to employ data-sparse approximations, such as hierarchical or low-rank representations [5, 8, 10, 17, 22, 23]. These methods accelerate matrix-vector multiplications by compressing the underlying structure of the matrix. Under this framework, we modify the results from Theorem 3.8 to get an error bound for approximating the square root of a preprocessed matrix using the Arnoldi iteration. This is summarized in the following theorem.
Theorem 3.13.
Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, assume that M admits a data-sparse approximation $\widetilde{M}$ which is positive definite and satisfies, for some $0<\epsilon<1/\kappa(M)$, $\|M-\widetilde{M}\|_{*}\leq\epsilon\|M\|$. Then, the relative error of the Arnoldi process for approximating the matrix square root is bounded by
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\epsilon^{1/2}\kappa^{1/2}(M)+\frac{4\sqrt{2}(1+\epsilon)^{7/2}(\kappa(M))^{9/2}}{(k-1/2)^{3/4}(1-\epsilon\kappa(M))^{9/2}}.$
Proof 3.14.
Based on $\widetilde{M}$, we apply the Arnoldi process to compute the matrix square root. Now the problem turns into:
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|}{\|M^{1/2}\bm{b}\|}+\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}.$
For the first term on the right hand side, according to the Powers–Størmer inequality [32], we have
$\frac{\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M^{1/2}-\widetilde{M}^{1/2}\|\|\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M-\widetilde{M}\|_{*}^{1/2}\|\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\epsilon^{1/2}\kappa^{1/2}(M).$
With Theorem 3.8, the second term is bounded by:
$\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|\widetilde{M}^{1/2}\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(\widetilde{M}))^{7/2}}{(k-1/2)^{3/4}}.$
By Weyl’s inequality [4], we have $|\sigma_{k}(M)-\sigma_{k}(\widetilde{M})|\leq\|M-\widetilde{M}\|\leq\|M-\widetilde{M}\|_{*}\leq\epsilon\|M\|=\epsilon\sigma_{\max}(M)$. Rearranging terms gives
$\kappa(\widetilde{M})\leq\frac{(1+\epsilon)\sigma_{\max}(M)}{\sigma_{\min}(M)-\epsilon\sigma_{\max}(M)}=\frac{(1+\epsilon)\kappa(M)}{1-\epsilon\kappa(M)}.$
Therefore, we conclude that
$\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}$ $\leq\frac{4\sqrt{2}(\kappa(\widetilde{M}))^{7/2}}{(k-1/2)^{3/4}}\frac{\|M^{1/2}\bm{b}\|}{\|\widetilde{M}^{1/2}\bm{b}\|}$
$\leq\frac{4\sqrt{2}(1+\epsilon)^{7/2}(\kappa(\widetilde{M}))^{9/2}}{(k-1/2)^{3/4}(1-\epsilon\kappa(M))^{9/2}}.$
@@ 4 Numerical experiments and discussion
In this section, we report on numerical studies of the Arnoldi iteration for computing the matrix square root applied to a vector, that is, $M^{1/2}\bm{b}$. First, we explore various matrix settings to study the dependence of relative error on the condition number of matrices and the iteration count. We start with a Hermitian matrix, confirming the result in Remark 3.10 as well as Lemma 3.3 in [1]. Then, we test both sparse and general non-Hermitian matrices validating Theorem 3.8. Finally, we explore the use of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ in the simulation of particulate suspensions to demonstrate the effectiveness of the proposed method in practical applications. Consistently across examples, the Arnoldi approximation is terminated based on the relative residual norm for solving $M\bm{x}=\bm{b}$, with a tolerance of $10^{-2}$. In particular, all residuals and relative errors reported in this section are measured in the 2-norm.
4.1 2D Laplacian
In this example, we consider the SPD matrix $M$ arising from a five-point stencil discretization of the Laplace equation with homogeneous Dirichlet boundary conditions on the unit square domain. The discrete problem is formulated on a uniform numerical grid with grid spacing $h=\frac{1}{n}$, where $n$ denotes the number of divisions along each axis. Then, $M$ takes the form of a block-structured matrix:
$M:=\begin{pmatrix}C&D&0&\cdots&0\\ D&C&D&\cdots&0\\ 0&D&C&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&D&C\end{pmatrix}$
where, $C$ and $D$ are defined as:
$C=\operatorname{tridiag}\left(-\frac{1}{h^{2}},\;\frac{4}{h^{2}},\;-\frac{1}{h^{2}}\right)_{(n-1)\times(n-1)},$
$D=\operatorname{diag}\left(-\frac{1}{h^{2}},-\frac{1}{h^{2}},\dots,-\frac{1}{h^{2}}\right)_{(n-1)\times(n-1)}.$
In this case, vector $\bm{b}$ is taken to be the vector of all ones. The performance of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ is presented in Table 1. The numerical experiments confirm that as the condition number of the matrix increases, maintaining a relative error on the order of $10^{-5}$ requires a larger number of iterations. The decay of the normalized residual for FOM and the relative error for the Arnoldi method are depicted in Fig. 2. In this example, a normalized residual of $10^{-2}$ for FOM yields a relative error of approximately $10^{-5}$ in the Arnoldi approximation, showing a three order of magnitude improvement. This observation indicates that employing the normalized residual of FOM as a stopping criterion is both effective and efficient, enabling accurate results even with a relatively large tolerance. To gain further insight into this example, we study the relationship between the relative error, condition number, and iteration count. The log-log plot of the relative error decay against the iteration count $k$ for different condition numbers is shown in Fig. 3. We observe that the error exhibits a decay rate of approximately $\mathcal{O}(k^{-3/2})$ with respect to the iteration count $k$. Additionally, we fix $k=25$ and plot the relative error against the condition number on a log-log scale in Fig. 4. As shown, the error decay scales with the condition number $\kappa$ at a rate close to $\mathcal{O}(\kappa^{5/2})$. These results are consistent with the conclusions for Hermitian matrices stated in Remark 3.10 and Lemma 3.3 in [1].
Table 1: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for the discrete 2D Laplacian.
$n$ $\kappa(M)$ Relative Error Iteration Counts $k$
30 529.97 1.90e-05 29
40 942.53 1.59e-05 39
50 1472.96 1.07e-05 50
60 2121.27 9.98e-06 60
70 2887.45 7.84e-06 71
80 3771.51 7.57e-06 81
90 4773.44 6.31e-06 92
100 5893.24 6.22e-06 102
110 7130.92 4.69e-06 114
Figure 2: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for the discrete 2D Laplacian.
Figure 3: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for the discrete 2D Laplacian.
Figure 4: Log-log plot of relative error with $k=25$ against condition number for the discrete 2D Laplacian.
4.2 Convection-Diffusion problem
In this case, $M$ is taken from the discretization of the convection-diffusion problem, which describes transport phenomena in a given domain. The continuous convection-diffusion operator is defined as
$Lu:=\eta u^{{\prime}{\prime}}(x)+u^{\prime}(x),\quad x\in[0,1],$

### v3

@@ 3.2 Matrix square root
If the matrix $M$ is Hermitian, then using the results from [chen2022error] and [amsel2023near], we obtain
$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$
$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\frac{\lambda_{\max}(M)}{\lambda_{\max}(M)+x}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\left(\frac{\lambda_{\max}(M)}{\lambda_{\max}(M)+x}\right)^{k+1}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
(12) $\leq\frac{1}{2\sqrt{\pi}}\frac{2k}{2k-1}(\lambda_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{2}}\|\bm{\xi}^{k}_{0}\|.$
This sharper result stems from the fact that the Hermitian matrix is unitarily diagonalizable with real eigenvalues, which allows for a more refined spectral analysis of $\left|\frac{\det(H_{k})}{\det(H_{k}-zI)}\right|$ and $\|(M+xI)^{-1}M\|$.
Thus, as a byproduct of our analysis, we derive an error bound for the Hermitian case that is tighter than (12), the result from [chen2022error].
Theorem 3.15.
Given a positive-definite Hermitian matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, the error of the Arnoldi process for approximating the matrix square root is bounded by
(13) $\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|\leq\frac{1}{2\sqrt{\pi}}\frac{2k}{2k-1}\left(\bar{\lambda}_{k+1}(M)\right)^{\frac{3}{2}}k^{-\frac{3}{2}}\left\|\bm{\xi}_{0}^{k}\right\|,$
where
$\bar{\lambda}_{k+1}(M):=\frac{1}{k+1}\left(\sum_{i=1}^{k}\lambda_{i}(M)+\lambda_{1}(M)\right)\leq\lambda_{\max}(M).$
Proof 3.16.
Let $\lambda_{1}(M)\geq\lambda_{2}(M)\geq\cdots\geq\lambda_{n}(M)$ be the eigenvalues of $M$. Then, $\lambda_{\max}(M):=\lambda_{1}(M)$. Using the arithmetic–geometric mean inequality, we obtain
$\prod_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\frac{\lambda_{1}(M)}{\lambda_{1}(M)+x}\leq\left(\frac{1}{k+1}\left(\sum_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}+\frac{\lambda_{1}(M)}{\lambda_{1}(M)+x}\right)\right)^{k+1}.$
Observing that $g_{x}(\lambda):=\frac{\lambda}{\lambda+x}$ is a concave function in $\lambda$, Jensen’s inequality yields
$\frac{1}{k+1}\left(\sum_{i=1}^{k}g_{x}(\lambda_{i}(H_{k}))+g_{x}(\lambda_{1}(M))\right)\leq g_{x}\left(\frac{\sum_{i=1}^{k}\lambda_{i}(H_{k})+\lambda_{1}(M)}{k+1}\right).$
Then,
$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$
$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\frac{\lambda_{1}(M)}{\lambda_{1}(M)+x}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\left(\frac{\frac{1}{k+1}\left(\sum_{i=1}^{k}\lambda_{i}(H_{k})+\lambda_{1}(M)\right)}{\frac{1}{k+1}\left(\sum_{i=1}^{k}\lambda_{i}(H_{k})+\lambda_{1}(M)\right)+x}\right)^{k+1}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Utilizing the interlacing inequality gives
$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$ $\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\left(\frac{\bar{\lambda}_{k+1}(M)}{\bar{\lambda}_{k+1}(M)+x}\right)^{k+1}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$\leq\frac{1}{2\sqrt{\pi}}\frac{2k}{2k-1}\left(\bar{\lambda}_{k+1}(M)\right)^{\frac{3}{2}}k^{-\frac{3}{2}}\left\|\bm{\xi}_{0}^{k}\right\|.$
This implies the bound (13) is sharper then the bound (12) almost for all cases.
Remark 3.17.
Additionally, one can similarly derive an error estimate for the Arnoldi approximation of the matrix inverse square root:
$\|M^{-1/2}\bm{b}-\operatorname{Arn}_{k}(x^{-1/2};M,\bm{b})\|\leq\frac{\Gamma(1/4)}{2^{3/4}\pi}\frac{2k}{2k-1}\left(\sigma_{\max}(M)\right)^{\frac{1}{2}}k^{-\frac{1}{4}}\|\bm{\xi}^{k}_{0}\|.$
If $M$ is Hermitian, then following [amsel2023near] and the discussion above, we can derive the following tighter bound, which also improves upon the state-of-the-art result,
$\|M^{-1/2}\bm{b}-\operatorname{Arn}_{k}(x^{-1/2};M,\bm{b})\|\leq\frac{1}{\sqrt{\pi}}\frac{2k+2}{2k+1}\left(\bar{\lambda}_{k+1}(M)\right)^{\frac{1}{2}}(k+1)^{-\frac{1}{2}}\|\bm{\xi}^{k}_{0}\|.$
@@ 3.3 Preprocessing
In practice, matrices arising from large-scale problems are often too large to be directly measured, stored, or analyzed, necessitating efficient preprocessing. To further reduce the cost, one direction is to employ data-sparse approximations, such as hierarchical or low-rank representations [bjarkason2019pass, clarkson2017low, datta2010numerical, hackbusch2015hierarchical, kishore2017literature, lin2011fast]. These methods accelerate matrix-vector multiplications by compressing the underlying structure of the matrix. Thus, we generalize the results from Theorem 3.8 to get an error bound for approximating the square root of a perturbed matrix using the Arnoldi iteration.
Corollary 3.18.
Let $M\in\mathbb{C}^{n\times n}$ be a positive-definite matrix such that for any unit vector $\bm{x}$, $\operatorname{Re}(\bm{x}^{*}M\bm{x})\geq\mu_{1}^{2}$ for some $\mu_{1}>0$, and let $\bm{b}\in\mathbb{C}^{n}$ be a given vector. We consider a perturbed matrix $\widetilde{M}$, satisfying $\operatorname{Re}(\bm{x}^{*}\widetilde{M}\bm{x})\geq\mu_{2}^{2}$ for some $\mu_{2}>0$, as well as $\|M-\widetilde{M}\|\leq\epsilon\|M\|$ for some $\epsilon>0$. Then, the error of the Arnoldi process for approximating the matrix square root is bounded by
$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|$
(14) $\leq\frac{\epsilon\sigma_{\max}(M)}{\mu_{1}+\mu_{2}}\|\bm{b}\|+\frac{\Gamma\!(3/4)(1+\epsilon)^{3/2}}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
Proof 3.19.
Based on $\widetilde{M}$, we apply the Arnoldi process to compute the matrix square root. Now the problem becomes
$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|\leq\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|+\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|.$
The first term can be bounded using the perturbation estimate in [schmitt1992perturbation], thus,
$\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|\leq\|M^{1/2}-\widetilde{M}^{1/2}\|\|\bm{b}\|\leq\frac{1}{\mu_{1}+\mu_{2}}\|M-\widetilde{M}\|\|\bm{b}\|\leq\frac{\epsilon\sigma_{\max}(M)}{\mu_{1}+\mu_{2}}\|\bm{b}\|.$
By Theorem 3.11, the second term is bounded by:
$\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|\leq\frac{\Gamma\!(3/4)}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(\widetilde{M}))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
By Weyl’s inequality [bhatia1997matrix], we have $|\sigma_{k}(M)-\sigma_{k}(\widetilde{M})|\leq\|M-\widetilde{M}\|\leq\epsilon\|M\|=\epsilon\sigma_{\max}(M)$. Combining the above estimates yields (3.18)
@@ 4 Numerical experiments and discussion
In this section, we report on numerical studies of the Arnoldi iteration for computing the matrix square root applied to a vector, that is, $M^{1/2}\bm{b}$. We first evaluate the a posteriori bounds (9) and (10), and the a priori bound (11) using non-Hermitian matrices. Next, we compare the a priori bounds (12) and (13) in the Hermitian setting to demonstrate the improvement yielded by our new bound (13). Then, we investigate the dependence of the approximation error on the largest singular value $\sigma_{\max}(M)$ and the iteration count $k$. Finally, we explore the use of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ in the simulation of particulate suspensions to demonstrate the effectiveness of the proposed method in practical applications. Consistently across examples, the Arnoldi approximation is terminated based on the relative residual norm for solving $M\bm{x}=\bm{b}$, with a tolerance of $10^{-2}$. In particular, all integrals in the bounds are evaluated numerically using integral in MATLAB with a relative tolerance of $10^{-8}$ and an absolute tolerance of $10^{-12}$.
4.1 Error bounds
Here, we compare the error bound estimates for cases when the spectrum of the matrix is uniformly distributed versus when the eigenvalues are clustered. We test this on both non-Hermitian and Hermitian examples.
4.1.1 The non-Hermitian case
Consider a positive definite non-Hermitian matrix $M\in\mathbb{R}^{500\times 500}$, constructed as $M=M_{0}+K$. The symmetric component $M_{0}$ is defined using $M_{0}=Q\mathrm{diag}(\Lambda)Q^{T}$, where $Q$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix. The matrix $K$ is introduced as a skew-symmetric term derived from a general random matrix $R$:
$K=\tfrac{1}{2}(R-R^{T}).$
First, we construct $\Lambda$ such that its eigenvalues are uniformly distributed in the interval $[1,1000]$. Fig. 2 displays the eigenvalues in the complex plane and the distribution of their modulus. We see that the eigenvalues of $M$ are widely distributed.
Figure 2: Eigenvalues of the uniformly-distributed spectrum matrix in the complex plane (left) and their magnitudes (right).
Setting $\bm{b}$ as an all-ones vector, the error of the Arnoldi process for computing $M^{1/2}\bm{b}$ are shown in Fig. 3. We observe that the intermediate bound Eq. 10 remains relatively sharp. The overestimation in the a priori bound Eq. 11 mainly arises from the rough estimate $|\lambda_{i}(H_{k})|\leq\sigma_{\max}(M)$, which is consistent with our earlier discussion in Remark 3.13.
Figure 3: Error of the Arnoldi process for a non-Hermitian matrix with a uniformly-distributed spectrum.
Next, let $\Lambda=[\lambda_{\text{cluster}};\lambda_{\text{outlier}}]$, where 95% of the eigenvalues are normally distributed around 1000:
$\lambda_{\text{cluster}}\sim 1000+100\,\mathcal{N}(0,1).$
The remaining outliers are drawn from a smaller normal distribution centered near 10:
$\lambda_{\text{outlier}}\sim 10+5\,\mathcal{N}(0,1).$
The distribution of eigenvalues and the magnitudes $|\lambda_{i}(M)|$ are displayed in Fig. 4. As shown, most of the $|\lambda_{i}(M)|$ lie around 1000, with only a few small isolated values.
Figure 4: Eigenvalues of the clustered spectrum matrix in the complex plane (left) and their magnitudes (right).
The vector $\bm{b}$ is taken to be the average of 100 eigenvectors corresponding to the 100 eigenvalues of largest magnitude. Then, the corresponding results of the Arnoldi process are presented in Fig. 5. Due to the clustered eigenvalue distribution, the method exhibits fast convergence, with a priori bounds (11) providing a sharper estimate in this regime as expected.
Figure 5: Error of the Arnoldi process for a non-Hermitian matrix with a clustered spectrum.
4.1.2 The Hermitian case
Next, we construct a $500\times 500$ Hermitian matrix as $M=Q\,\mathrm{diag}(\Lambda)\,Q^{T}$, where the eigenvalues of the diagonal matrix $\Lambda$ follow a uniform distribution over $[1,1000]$. In Fig. 6, we report the error and bounds in Remark 3.14 and Theorem 3.15. We observe that, as $k$ increases, the Ritz value $\lambda_{i}(H_{k})$ approaches a true eigenvalue of $M$, leading the bound (9) to become sharper and align more closely with the actual error. We also plot the evolution of $\bar{\lambda}_{k+1}(M)$, with $\sigma_{\max}(M)$ shown in Fig. 7 for comparison. It can be observed that when the eigenvalues are uniformly distributed and lie close to one another, bound (13) yields only a minimal improvement over bound (12).
Figure 6: Error of the Arnoldi process for a Hermitian matrix with a uniformly-distributed spectrum.
Figure 7: Evolution of $\bar{\lambda}_{k+1}$ for a symmetric matrix with a uniformly-distributed spectrum.
Next, we take $\Lambda=[\lambda_{\text{cluster}};\,\lambda_{\text{outlier}}]$ with 99% of the eigenvalues tightly clustered near 10, and the remaining 1% sampled as outliers around 1000. The same $\bm{b}$ is used. The results are displayed in Fig. 8 and Fig. 9. Clearly, as $k$ increases, the sequence $\bar{\lambda}_{k+1}(M)$ decreases monotonically, and becomes progressively smaller than $\sigma_{\max}(M)$. Therefore, as expected in Remark 3.13, when the larger eigenvalues are more widely spread, bound (13) exhibits a tighter estimate compared to bound (12).
Figure 8: Error of the Arnoldi process for a symmetric matrix with a clustered spectrum.
Figure 9: Evolution of $\bar{\lambda}_{k+1}$ for a symmetric matrix with a clustered spectrum.
4.2 Scaling behavior
In this example, we test the algorithm’s dependence on the number of Arnoldi iterations $k$ and the largest singular value $\sigma_{\max}(M)$. Here, we construct $M$ from the discretization of the convection-diffusion problem. The continuous convection-diffusion operator is defined as
$Lu:=-\eta u^{{\prime}{\prime}}(x)+u^{\prime}(x),\quad x\in[0,1],$